A golden rectangle is a rectangle whose length is approximately 1.6 times its width. The early Greeks thought that a rectangle with these dimensions was the most pleasing to the eye and examples of the golden rectangle are found in many early works of art. For​ example, the Parthenon in Athens contains many examples of golden rectangles. Mike Hallahan would like to plant a rectangular garden in the shape of a golden rectangle. If he has 234234 feet of fencing​ available, find the dimensions of the garden.

Answer by ankor@dixie-net.com(22740)

A golden rectangle is a rectangle whose length is approximately 1.6 times its width.
Mike Hallahan would like to plant a rectangular garden in the shape of a golden rectangle.
If he has "234"* feet of fencing is available, find the dimensions of the garden. *Makes more sense.
:
let w = the width of the garden
then
1.6w = the length
:
2(1.6w) + 2w = 234
3.2w + 2w = 234
5.2w = 234
w = 234/5.2
w = 45 ft is the width
and
1.6(45) = 72 ft is the length
:
:
Check
2(72) + 2(45) =
144 + 90 = 234